If a normal population of overpayments by National Health Insurance Authority to a private health service


provider is known to have a population standard deviation equal to GHC5, how large a sample would we


take in order to be 95% confident that the sample mean overpayment will not differ from the population


mean overpayment by more than ±0.80?

1
Expert's answer
2022-03-19T04:09:11-0400

"n=( \frac{Z_{\alpha/2}}{\sigma})^2p(1-p)"

"Z_{\alpha/2}(0.95)=1.65"

"n=(1.65\/0.05)^20.8(1-0.8)=174"